Calculate S_35 for the arithmetic sequence in which a_5=19 and the common difference is d=-1.3.
Answer by rothauserc(4718)

the formula for the nth term in an arithmetic sequence is
:
a(n) = a(1) + d(n-1)
:
we are given
:
a(5) = 19 = a(1) -1.3(5-1)
a(1) = 19 +5.2 = 24.2
:
the formula for the sum of the first nth terms of an arithmetic sequence is
:
S(n) = (n/2)(a(1) + a(n))
:
a(35) = 24.2 -1.3(35-1) = -20
:
********************************
S(35) = (35/2)(24.2 -20) = 73.5
********************************
: